The surface area of a cube is 50 cm squared.
what is the area of one face of the cube? What is the length of one edge of the cube? How do you find the answer?

Answer by stanbon(75887)

The 6 faces of the cube all have the same area.
Let a side of a face be x; then its area is x^2
EQUATION: 6x^2=50
x^2=25/3 sq. cm. (Area of one face)
x=5/sqrt3 = 5sqrt3/3= (5/3)sqrt3 (length of one side)
Cheers,
Stan H.